OpenStudy (vishweshshrimali5):

A positive number when divided by 88 gives the remainder 8. What will be the remainder when this number id divided by 11 ?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

lets try it. what number when divided by 88 gives a remainder of 8? well 88+8=96 for for one

OpenStudy (vishweshshrimali5):

Yes I have done hit and trial and what i have got is that the remainder = 8

OpenStudy (anonymous):

so i guess 3 is wrong because \[\frac{96}{11}=8\tfrac{8}{11}\] so remainder is 8. shows you what i know!

OpenStudy (vishweshshrimali5):

But how will you prove it

OpenStudy (vishweshshrimali5):

I want the whole explanation for it

OpenStudy (anonymous):

and we can try it again with \[2\times 88 + 8\]

OpenStudy (anonymous):

oh now that we know it is 8, i think answer is easy enough

OpenStudy (vishweshshrimali5):

Yes that also gives remainder 8

OpenStudy (vishweshshrimali5):

But how will you prove it

OpenStudy (anonymous):

because \[88=8\times 11\] yes?

OpenStudy (vishweshshrimali5):

Right !

OpenStudy (vishweshshrimali5):

Now what to do ?

OpenStudy (anonymous):

so \[n\times 88+8=11n+8\]

OpenStudy (vishweshshrimali5):

Now !

OpenStudy (anonymous):

and when you divide \[11n+8\] by 11 you get a remainder of 8

OpenStudy (vishweshshrimali5):

How ?

OpenStudy (anonymous):

since clearly11 divides 11n evenly!

OpenStudy (vishweshshrimali5):

Yes

OpenStudy (vishweshshrimali5):

And we get 8 as remainder

OpenStudy (vishweshshrimali5):

Thanks

OpenStudy (anonymous):

that is all you need. 11 goes in to "11n" n times, so it goes in to 11n + 8 n times with a remainder of 8.yw

OpenStudy (vishweshshrimali5):

Thanks

OpenStudy (vishweshshrimali5):

But how did you get n×88+8=11n+8

OpenStudy (vishweshshrimali5):

I can't get that still ?

OpenStudy (vishweshshrimali5):

Because if I will show that using hit and trial I got this My teacher will give me zero marks

OpenStudy (vishweshshrimali5):

Please help !!!!

OpenStudy (anonymous):

what to you mean "hit and trial"? we proved it

OpenStudy (vishweshshrimali5):

How did we prove that n×88+8=11n+8

OpenStudy (anonymous):

just say "if n is divided by 88 and the remainder is 8, then \[n=88k+8\] and therefore \[n=11\times 8j+8\] since \[88=11\times 8\]

OpenStudy (vishweshshrimali5):

Ok Then ?

OpenStudy (anonymous):

'thus upon dividing \[n=11\times 8j+8\] by 11 we see that 11 divides \[11\times 8j\] evenly giving a remainder of 8" qed aka have a nice day

OpenStudy (vishweshshrimali5):

Thanks a lot satellite

OpenStudy (vishweshshrimali5):

Will you please help me in some more questions

OpenStudy (vishweshshrimali5):

Go to my profile page and then to the questions asked

OpenStudy (anonymous):

post them be happy to

OpenStudy (vishweshshrimali5):

Ok